OpenStudy (anonymous):

A hamster ball has a radius of 18 centimeters. What is its volume written in terms of π? Answer 432π cm3 4,374π cm3 5,832π cm3 7,776π cm3

OpenStudy (anonymous):

i got 324 but thats not right

OpenStudy (ash2326):

Hamster ball is a sphere. A sphere's volume is given as \[\frac{4}{3} \pi r^3\] r=18 cm so \[ \frac{4}{3} \pi (18)^3=\frac{4}{3} \pi 324 \times 18=> \pi \times 4 \times 108 \times 18\] we get \[ \pi7776\]